OpenStudy (anonymous):

certain regions is losing 12 million acres if rain forest each year a)find a linear function f that calculates the change in acres of rain forest in million in x years b) find f(4) and interpret the results a) f(x)= b f(4)= after ?? years the regions can expect to have lost about ?? million of acres of rain forest

OpenStudy (anonymous):

It's saying that the rainforest loses ("changes") 12 million acres/year. amount_lost or amount_changed = 12 * years

OpenStudy (anonymous):

So, f(years) = 12 * years

OpenStudy (anonymous):

so a) f(x)=12 and b) f(4)= 48

OpenStudy (anonymous):

so in 48 years they region can lost about 576 million acres

OpenStudy (anonymous):

If f(x) = 12 Then how do you get f(4) = 48. Show your calculations. (Hint: you can't.)

OpenStudy (anonymous):

12*4

OpenStudy (anonymous):

In your equation f(x) = 12, there's no place to put the 4.

OpenStudy (anonymous):

ok but the question is saying for part b f(4)

OpenStudy (anonymous):

Yep, but for part a the answer is: f(x) = 12 * x

OpenStudy (anonymous):

so 12*12

OpenStudy (anonymous):

Nope. x is the number of years. f(x) = 12 * x So f(4) = 12 * 4 =48

OpenStudy (anonymous):

so the answer would be in 48 years the region can expect 576 million acres

OpenStudy (anonymous):

why is this math so hard I am dying here, been trying to do this exam for 9 hors now

OpenStudy (anonymous):

You're trying to find the number of acres in 4 years, right? (Not 48?)

OpenStudy (anonymous):

Try to keep organized as you think about this. First figure out what x means -- it's the number of years. Then figure out what f(x) means -- it's the amount of acres lost (over x years).

OpenStudy (anonymous):

that right so it would be 4 so in 4 years they would lose 48 million acres right

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

so this is what my answer would look like a) f(x)=12 b) f(4)= 48 after 4 years the rain forest lost 48 million acres

OpenStudy (anonymous):

a) is actually 12*x Your answer should be: a) f(x) =12x b) f(4) = 48 After 4 years the rain forest lost 48 million acres.